Show that the average energy of a weakly damped oscillator is given by:
< E > = E0 exp (-2bt)
1
Expert's answer
2020-02-27T09:58:23-0500

The energy of the system is


"E(t)=KE(t)+U(t)=\frac{1}{2}m\bigg(\frac{dx}{dt}\bigg)^2+\frac{1}{2}kx^2."

The instantaneous displacement of a weakly damped harmonic oscillator is


"x (t) = a_0 \text{exp} (- bt)\cdot \text{cos} (\omega_dt + \phi),\\
\space\\
\frac{dx(t)}{dt}= - a_0 \text{exp} (- bt) [b \text{cos} (\omega_dt + \phi) + \omega_d\text{sin} (\omega_dt + \phi)],\\
\space\\
KE(t)=\frac{1}{2}ma_0^2 \text{exp}(-2bt) [b\text{cos} (\omega_dt + \phi) + \omega_d\text{sin} (\omega_dt + \phi)]^2"


(open the squared parentheses).

The potential energy is


"U=\frac{1}{2}kx^2=\\
\space\\
=\frac{1}{2}ka_0^2 \text{exp} (-2bt)\cdot \text{cos}^2 (\omega_dt + \phi)."

Remember that


"k=m\omega_0^2."

"U=\frac{1}{2}kx^2=\\
\space\\
=\frac{1}{2}m\omega_0^2a_0^2 \text{exp} (-2bt)\cdot \text{cos}^2 (\omega_dt + \phi)."

During one oscillation, for small amplitudes, the average values of sine and cosine squared are equal


"\text{sin}^2 (\omega_dt + \phi)=\text{cos}^2 (\omega_dt + \phi)=\frac{1}{2}."

Remember that the average values of sine per one oscillation is 0.

Therefore, after adding the KE with opened parentheses and U, we get


"<E>=\frac{1}{2}ma_0^2 \text{exp}(-2bt)\bigg[\frac{b^2+\omega_0^2}{2}+\frac{\omega_d^2}{2}\bigg]=\\
\space\\
=\frac{1}{2}ma_0^2\omega^2_0\text{exp}(-2bt)."

Since


"\frac{1}{2}ma_0^2\omega^2_0=E_0,"

we can finally write

"<E>=E_0\text{exp}(-2bt)."